# Algebraic Robustness Verification of Neural Networks

### v1

@@ B Proofs
By Lemma˜B.2, the joint distribution of $(f(\bm{x}),\nabla f(\bm{x}))$ is Gaussian with mean zero and covariance
@@
where the last equality follows from the block matrix inversion formula via Schur complement. To find $\mathcal{I}(\bm{x},\lambda)$, we need to find the conditional distribution of $\nabla^{2}f(\bm{x})$ given $f(\bm{x})=0$ and $\nabla f(\bm{x})=\frac{\bm{\xi}-\bm{x}}{\lambda}$. Since the kernel $K_{f}$ is $C^{4}$, by a similar argument as in Lemma˜B.2, the joint distribution of $\left(f(\bm{x}),\nabla f(\bm{x}),\operatorname{vech}{(\nabla^{2}f(\bm{x}))}\right)$ is still Gaussian. The covariance matrix can be computed using the derivatives of the kernel $K_{f}$, i.e.,
@@
Lemma B.2.
@@
B.7 Proof of Corollary˜3.11
See 3.11
@@
When the input dimension of $f$ is $n=1$, the hypersurface $\{x\in\mathbb{R}:f(x)=0\}$ degenerates to a finite set of points. In this the case, the number of real ED critical points coincides with the number of real zeros of the random function $f(x)=0$. By Proposition˜B.3, $f$ is a centered Gaussian process with kernel $K(x,x^{\prime})$ defined as
$K(x,x)=\operatorname{Var}[f(x)]=(2d-1)!!(x^{2}+1)^{d}.$
We define the Kac–Rice density as
$\rho(x)$ $=\int_{\mathbb{R}}|y|\,p_{(f(x),f^{\prime}(x))}(0,y)\,dy$
$=\int_{\mathbb{R}}|y|\,p_{(f(x),f^{\prime}(x))}(0,y)\,dy.$
Then we have
$=\frac{1}{2\pi\sqrt{\det K_{F}(x,x)}}\cdot\frac{2\det K_{F}(x,x)}{K(x,x)}$
$=\frac{\sqrt{\det K_{F}(x,x)}}{\pi K(x,x)}.$
By Proposition˜B.5, we have
The determinant of $K_{F}(x,x)$ can be computed as
$\quad-\left[d(2d-1)!!\,x(x^{2}+1)^{d-1}\right]^{2}$
$=d^{2}(2d-1){((2d-3)!!)}^{2}(x^{2}+1)^{2d-2}.$
The Kac–Rice density can then be found by
Fig.˜5 shows a simulation study that verifies the above theoretical result.
Proposition B.3 (NNGP of shallow polynomial networks).
@@
Remark B.4.
We provide two identities about the coefficients $c(s,d)$ appearing in Eq.˜9. These identites can be useful in simplifying certain expressions involving the kernel.
@@
Proposition B.5.
Given the kernel function $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ defined as in Eq.˜9, its first and second derivatives evaluated at the same point $\bm{x}^{\prime}=\bm{x}$ are given by
@@
B.8 Proof of Proposition˜4.5
@@
C Experiments and Implementation Details
Benchmark configurations
Data generation
  • Set $\delta=m\cdot v$.
Model architecture and training
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{margin}}$
where $f_{k}(x)$ denotes the logit for class $k$, and $m=0.01$ is the margin.
where $T=5,000$ is the total number of epochs.
Training Settings Results
6 2 0.2 100% 60% 0.0448(0.0063)0.0448\,(0.0063)
6 2 0.5 100% 60% 0.0546(0.0242)0.0546\,(0.0242)
6 3 0.2 100% 20% 5.8013(0.5377)5.8013\,(0.5377)
6 3 0.5 100% 40% 5.4401(0.2486)5.4401\,(0.2486)
10 2 0.2 100% 30% 0.0331(0.0063)0.0331\,(0.0063)
10 2 0.5 100% 50% 0.0335(0.0102)0.0335\,(0.0102)
10 3 0.2 100% 30% 4.6127(0.1721)4.6127\,(0.1721)
10 3 0.5 100% 50% 4.4646(0.2993)4.4646\,(0.2993)
Verification
For instances designated as unverifiable by construction, a sound verification method must always report falsification, since counterexamples are guaranteed to exist. Any verification claim on an unverifiable instance therefore constitutes a false positive and signals a soundness error in the verifier. In contrast, for clean instances, both verification and falsification are acceptable outcomes, depending on the true robustness of the model at the given input. This distinction allows us to separately evaluate verifier soundness on unverifiable instances and verifier precision on clean instances. Table˜2 summarizes the verification results across all 8 benchmark configurations. The verifier achieves perfect soundness on all unverifiable instances, correctly identifying all planted counterexamples. On clean instances, the verifier successfully falsifies a significant fraction, demonstrating its ability to detect non-robustness even without planted counterexamples. The average verification time per instance is also reported, showing efficient performance across configurations.

### v2

@@ F Auxiliary Lemmas and Proofs for Expected Real ED Degree
By Lemma˜F.2, the joint distribution of $(f(\bm{x}),\nabla f(\bm{x}))$ is Gaussian with mean zero and covariance
@@
where the last equality follows from the block matrix inversion formula via Schur complement. To find $\mathcal{I}(\bm{x},\lambda)$, we need to find the conditional distribution of $\nabla^{2}f(\bm{x})$ given $f(\bm{x})=0$ and $\nabla f(\bm{x})=\frac{\bm{\xi}-\bm{x}}{\lambda}$. Since the kernel $K_{f}$ is $C^{4}$, by a similar argument as in Lemma˜F.2, the joint distribution of $\left(f(\bm{x}),\nabla f(\bm{x}),\operatorname{vech}{(\nabla^{2}f(\bm{x}))}\right)$ is still Gaussian. The covariance matrix can be computed using the derivatives of the kernel $K_{f}$, i.e.,
@@
F.2 Joint Gaussianity of the Augmented Process
Lemma F.2.
@@ F.2 Joint Gaussianity of the Augmented Process
F.3 Operator Norm Bound for the Adjoint Matrix
Lemma F.3.
For any $H\in\mathbb{R}^{n\times n}$ and $\lambda\in\mathbb{R}$,
$\|\operatorname{adj}(I_{n}+\lambda H)\|_{\mathrm{op}}\leq\bigl(1+|\lambda|\,\|H\|_{\mathrm{op}}\bigr)^{n-1}.$
@@ F.3 Operator Norm Bound for the Adjoint Matrix
Let $A:=I_{n}+\lambda H$, and let $\sigma_{1}(A)\geq\cdots\geq\sigma_{n}(A)\geq 0$ be its singular values. The singular values of $\operatorname{adj}(A)$ are
$\left\{\prod_{j\neq i}\sigma_{j}(A):\,i=1,\dots,n\right\},$
hence
$\|\operatorname{adj}(A)\|_{\mathrm{op}}=\max_{1\leq i\leq n}\prod_{j\neq i}\sigma_{j}(A)\leq\sigma_{1}(A)^{n-1}=\|A\|_{\mathrm{op}}^{\,n-1}.$
By the triangle inequality,
$\|A\|_{\mathrm{op}}=\|I_{n}+\lambda H\|_{\mathrm{op}}\leq\|I_{n}\|_{\mathrm{op}}+|\lambda|\,\|H\|_{\mathrm{op}}=1+|\lambda|\,\|H\|_{\mathrm{op}}.$
Combining the two inequalities gives
$\|\operatorname{adj}(I_{n}+\lambda H)\|_{\mathrm{op}}\leq\|I_{n}+\lambda H\|_{\mathrm{op}}^{\,n-1}\leq\bigl(1+|\lambda|\,\|H\|_{\mathrm{op}}\bigr)^{n-1}.$
F.4 NNGP of Shallow Polynomial Networks
Proposition F.4 (NNGP of shallow polynomial networks).
@@ F.4 NNGP of Shallow Polynomial Networks
Remark F.5.
We provide two identities about the coefficients $c(s,d)$ appearing in Eq.˜31. These identites can be useful in simplifying certain expressions involving the kernel.
@@
F.5 Derivatives of the NNGP Kernel
Proposition F.6.
Given the kernel function $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ defined as in Eq.˜31, its first and second derivatives evaluated at the same point $\bm{x}^{\prime}=\bm{x}$ are given by
@@ F.5 Derivatives of the NNGP Kernel
G Parameter Discriminant
G.1 Proof of Proposition˜4.5
@@ G Parameter Discriminant
H ReLU networks
Although our results primarily concern algebraic classifiers, the metric-projection formulation of Section 2 also gives a natural geometric interpretation of ReLU networks. By Proposition 2.1, which applies to arbitrary continuous classifiers, the robustness margin of a test point $\bm{\xi}$ with a unique predicted class may be computed by projecting $\bm{\xi}$ onto the relaxed boundary $\mathcal{V}_{c,c^{\prime}}^{\bm{\theta}}$. For a ReLU network, the input space $\mathbb{R}^{n}$ is partitioned into activation regions, and on each such region the map $f_{\bm{\theta}}$ is affine linear. Thus, for every full-dimensional activation region $P$ and every pair of classes $c\neq c^{\prime}$, there exist $a_{P,c,c^{\prime}}\in\mathbb{R}^{n}$ and $b_{P,c,c^{\prime}}\in\mathbb{R}$ such that
$f_{\bm{\theta},c^{\prime}}(\bm{x})-f_{\bm{\theta},c}(\bm{x})=a_{P,c,c^{\prime}}^{\top}\bm{x}+b_{P,c,c^{\prime}}\qquad\text{for all }\bm{x}\in P.$
Hence
$\mathcal{V}_{c,c^{\prime}}^{\bm{\theta}}\cap P=P\cap\{\bm{x}\in\mathbb{R}^{n}:a_{P,c,c^{\prime}}^{\top}\bm{x}+b_{P,c,c^{\prime}}=0\}.$
In particular, on each activation region, the relaxed boundary is an affine linear variety, and the corresponding local metric-projection problem has ED degree $1$.
Thus the ReLU case separates local and global complexity in a particularly transparent way. Locally, every piece of the relaxed boundary has ED degree 1, so there is no nontrivial algebraic complexity on a fixed activation region. Globally, however, the relaxed boundary is a union of such pieces across all activation regions, glued along lower-dimensional faces; the difficulty of verification is therefore governed by the combinatorics of this polyhedral stratification, namely by the number of relevant regions and how their boundary pieces fit together. For algebraic classifiers, by contrast, the relaxed boundary is already genuinely nonlinear on a single piece, and the ED degree measures this intrinsic local complexity.
This separation of local and global complexity also suggests the correct analog of the ED discriminant in the ReLU setting. On each fixed activation region, the relaxed boundary is an affine linear variety, so the local ED problem has a unique critical point and hence empty ED discriminant. Thus any change in the verification landscape must come from the global piecewise-linear geometry, not from local collisions of critical points: it occurs when the nearest point switches from one affine stratum to another, or moves onto a lower-dimensional face where strata meet. In this sense, the polyhedral separation of the activation regions plays, for ReLU networks, the same global role that the ED discriminant plays for algebraic decision boundaries. This also points to a natural intermediate class for future study, namely piecewise-algebraic classifiers such as networks with polynomial spline activations, where one expects both phenomena to coexist: nontrivial local ED geometry on each piece together with global complexity coming from how the pieces fit together.